40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 5 18:34:26 2009, Jenne, Update, Adaptive Filtering, Attempts to take a TF of the OAF system OAF_TF_CORRexc_EMPHout_2.png
    Reply  Tue Oct 6 11:13:53 2009, Jenne, Update, Adaptive Filtering, Attempts to take a TF of the OAF system 
       Reply  Wed Oct 7 03:49:49 2009, rana, Update, Adaptive Filtering, Attempts to take a TF of the OAF system Untitled.png
          Reply  Wed Oct 7 19:23:49 2009, Jenne, Update, Adaptive Filtering, (Final?) PEM cabling changes 
       Reply  Wed Oct 7 07:42:55 2009, rana, Update, Adaptive Filtering, Attempts to take a TF of the OAF system 
Message ID: 2061     Entry time: Wed Oct 7 03:49:49 2009     In reply to: 2058     Reply to this: 2065
Author: rana 
Type: Update 
Category: Adaptive Filtering 
Subject: Attempts to take a TF of the OAF system 

 Here's a plot of the spectra of the seismometers and MCL. The coherence shows which axes are aligned right now: MC1_X is coherent with GUR_NS which means that its mis-oriented.

I've now swapped the "MC1" cables: so the old "NS" now goes into EW and the old EW now goes into NS. VERT is unchanged.

Also fixed the channel names - the Guralp previously named MC1 is now GUR1 and the other one is GUR2. Also no more EW, NS, & VERT. Its all XYZ.

DAQD restarted with the new channel names.

Attachment 1: Untitled.png  69 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-